At 10:02AM, the N2 Pressure fell below 60 PSI. The watch script saw this happen, but I did not recieve the email it is supposed to send
C1:Vac-P1_pressure reads 7e-4, which is the same as it has for the past ~2 days, so the V1 interlock worked fine.
I've put some fresh N2 into the system, and Bob will pop in over the weekend to check it. I'll stay on top of it until Steve gets back.
After consulting ELOGs and the 40m wiki, I reasoned it was ok to open the V1 to reconnect the turbo pump to the main IFO volume and VM1 to reconnect the RGA, and have now done so.